OpenStudy (anonymous):

Solve -5(3n + 4) = 40. I somehow got -3 is that correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

start by distributing, what do you get?

OpenStudy (anonymous):

\[-5(3n+4)=40\]

OpenStudy (anonymous):

-5(3n)+(-5)(4) ?

OpenStudy (anonymous):

simplified 40= -15n-20 corect?

OpenStudy (anonymous):

distribute to get -5*3n + (-5*4)=40

OpenStudy (anonymous):

next get n by itself

OpenStudy (anonymous):

then simplify further to get -15n - 20=40. add 20 to both sides to get -15n=60 divide both sides by -15 to get n=-4

OpenStudy (anonymous):

ahaha i just did that too thanks, i got it.